Reports 1-1 of 1 Clear search Modify search
AdV-DAQ (Calibration)
rolland - 14:30 Thursday 20 November 2025 (68201) Print this report
Bad synchronisation of suspension DSPs

Following the recovery of this morning, there are alerts in the DMS (triggered only in LN3 mde): some DSPs from different suspensions are not properly synchronized:

  • the NE, WI and PR MIRrors have a different synchronisation that before the timing issue of last Monday. In particular the case of the NE mirror will introduce large errors in the reconstructed h(t) data.
  • the synchronisaton of the NE and WE MARionnettes is not activated since few months, so no information on them. But since they are not used for the controls, it has no impact on Hrec.

The first two VIM plots from today show the current status for the mirrors and for the marionettes  (the phase is computed at a line injected at 2012 Hz):

  • NE Mirror DSP off by 50 µs   --> impact on Hrec
  • WI Mirror DSP off by 50 µs   
  • PR Mirror DSP off by 50 µs   
  • SR MIrror DSP off by 25 µs 
  • WI Marionette DPS off by 25 µs   --> impact on Hrec

The figures 3 and 4 show the monitoring data on Monday 17 November: the synchronisation of the different DSPs changed at different times in the afternoon.

The last figure shows the current bias on hrec (supposed to be unbiased) : this has a huge impact on the phase bias,  (apparently not on the h(t) amplitude, which I do not undertand at first sight). The phase bias looks consistent with ~50 µs bias around 250 Hz.

At least the NE mirror and WI marionette DSPs should be resynchronized as soon as possible to have a correct reconstructed h(t).

 

Images attached to this report
Comments to this report:
rolland - 16:44 Friday 21 November 2025 (68216) Print this report

Yesterday afternoon and today, Valerio and the SAT team have reloaded and/or rebooted the different suspension DSPs to re-synchronize them.

As shown in the attached figures, the monitoring channels are back to their nominal values for all the suspensions (including SR).  Note that for online dataDisplay, the channels are available only late in the data collection pipeline, i.e. in the FbmStDy stream, but not in the other online streams. 

For NE and WE marionettes, there is no monitoring and the alert shelving in the DMS ended this week. I have shelved again the 6 associated flags(until January 12th 13h30).

 

 

Images attached to this comment
Search Help
×

Warning

Error

The present report has been modified outside this window. Please check for its integrity in the main page.

Refreshing this page will move this report into drafts.

×